OpenStudy (anonymous):

solve: √15y^2v^7√5yv^6

OpenStudy (anonymous):

\(\Large \sqrt{15y^2v^7} \sqrt{5yv^6} = \sqrt{(15y^2v^7)(5yv^6)}\\ \Large =\sqrt{75y^3v^{13}}=\sqrt{(25y^2v^{12})(3yv)}\\ \Large =5yv^6\sqrt{3yv}\)